the reaction between methane and water vapour produces carbon dioxide gas and hydrogen gas in the following reaction: CH4(g)+2H2O(g)=CO2(g)+4H2(g)

you initially add 2.00 mol of methane and 6.00 of water vapour to an 8.00L container. calculate Keq for the forward reaction if the equilibrium concentration of water vapour was measured to be 0.56mol/L


1
Expert's answer
2021-11-24T12:22:18-0500

Molar mass of methane=16.04= 2/16.04

= 0.125

Molar mass of H2O= 18

= 6/18= 0.33

= 0.455/8

= 0.0569mol/L